Write a system of two equations in two unknowns and solve by the method of your choice. Don’t forget to show the methodology. Irving has 35 coins, consisting of only dimes and nickels. If the value of his coins is $3.30, how many nickels and dimes does he have? Suppose, in the problem above, that Irving’s total cash was only $2.50. Then, how many nickels and dimes would he have? Again, solve by the method of your choice.
Answer by checkley77(12844)

d+n=35 or d=35-n
.05n+.10d=3.30
.05n+.10(35-n)=3.30
.05n+3.50-.10n=3.30
-.05n=3.30-3.50
-.05n=-.20
n=-.20/-.05
n=4 nickels.
d=35-4=31 dimes.
Proof:
.05*4+.10*31=3.30
.20+3.10=3.30
3.30=3.30